# Excluding cosmetic surgeries on hyperbolic 3-manifolds

### v2

@@ 4.2. Results
The cosmetic surgery conjecture, Conjecture 1.2, holds for the 59,107 one-cusped manifolds in the SnapPy census, provided that SnapPy’s systole calculation is correct to within 1%.
@@
For each manifold $N$ in the census, the procedure runs to completion. Each time, the set of unidentified slopes is $\mathcal{U}(N)=\emptyset$. (This is a consequence of Dunfield’s prior work [12, Theorem 1.2].) The procedure successfully distinguishes all exceptional pairs in $\mathcal{E}\times\mathcal{E}$ and all hyperbolic pairs in $\mathcal{H}^{p}\times\mathcal{T}^{p}$, apart from exactly 7 slope pairs that are described in detail below. In each case, the two fillings of $N$ produce oppositely oriented copies of the same closed 3-manifold $M$, where $M$ is chiral. Thus these fillings do not contradict Conjecture 1.2, which concerns purely cosmetic fillings.
@@
[…]
@@
[…]
@@
[…]
    This verification problem is addressed by taking covers. The double covers of $N(s)$ and $N(t)$ do admit positively oriented triangulations. By a theorem of Hilden, Lozano, and Montesinos [21, Theorem 2.2], the Chern–Simons invariant $\operatorname{CS}(M)$ is multiplicative under covers, modulo $\pi^{2}/2$. The manifold $M=N(s)$ has six distinct double covers up to isomorphism. Each double cover $\widehat{M}$ satisfies $\operatorname{CS}(\widehat{M})=-0.8224...\mod\pi^{2}/2$, which is verifiably distinct from 00. Thus $CS(M)=-0.4112...\mod\pi^{2}/4$, which is again verifiably distinct from 00. Since $CS(N(s))\neq 0$, we know that $N(s)$ is chiral, hence cannot admit an orientation-preserving isometry to $N(t)$.
[…]
@@
Since each of the above slope pairs produce distinct oriented 3-manifolds, we conclude that Conjecture 1.2 holds for every $N$ in the census. Recall from Remark 4.3 that this verification relies on the correctness of systole computations, up to a numerical error of less than $1\%$. ∎
@@
When ${\operatorname{sys}}(N)$ is very short, the program has to work much harder, for two reasons. First, the Dirichlet domain of such a manifold is very elongated due to the presence of a deep Margulis tube, making the domain (and hence the systole) hard to compute. Second, the set of short slopes $\mathcal{S}(N)$ computed in (4.1) has size approximately $2\pi\cdot{\operatorname{sys}}(N)^{-1}$. After filtering the hyperbolic slopes by homology, as in Steps 4 and 5, the number of slope pairs to check can be on the order of ${\operatorname{sys}}(N)^{-3/2}$. The example below shows that the first issue dwarfs the second.
[…]
Moving beyond this extreme example, the program encountered 294 one-cusped manifolds whose systole is shorter than $0.004$. Each of those systoles took over an hour to compute.
@@
[…]
@@
The chirally cosmetic surgery conjecture for knots in $S^{3}$, Conjecture 1.6, holds on the set of all hyperbolic knots up to 15 crossings, provided that SnapPy’s systole calculation is correct to within 1%.
@@
Theorem 4.7 was proved by running the python program check_mfds_chiral on the set of 313,209 hyperbolic knots up to 15 crossings. For most of these knots, the program takes only a few seconds to run to completion. However, there is a small list of 14–crossing knots, for which chirally cosmetic surgeries are particularly hard to distinguish:
@@
These 13 knots were previously identified by Stoimenow [40] as being almost amphicheiral: that is, each $K$ in the list is a mutant of its mirror image. Thus, for $N=S^{3}-K$, the Dehn filling $N(p,q)$ is a genus–$2$ mutant of the mirror image of $N(p,-q)$. Since complex hyperbolic volume is preserved under mutation [37, Theorem 1.1 and Corollary 1.2], it cannot distinguish $N(p,q)$ from $N(p,-q)$. In practice, we distinguished these fillings using the cover homology invariant $\mathcal{C}(N(s),7)$ described in Definition 3.1, using covers up to degree $7$.
The hardest knot to handle computationally turned out to be 14a12813. The computation of enough cover homology groups to distinguish its Dehn fillings took over 6 days of wall time making full use of 8 cores, for a total of 48 days and 9 hours of CPU time.
@@ 5. Common fillings of different parent manifolds
In this section, we describe how the procedure of Section 4 can be adapted to determine common Dehn fillings of two distinct manifolds. The results of this section were motivated by a question from Kalfagianni and Melby. In their joint work [32], they use our results (specifically Theorem 5.2) to find many new $q$–hyperbolic knots in $S^{3}$.
@@ 5.1. The procedure
Here, we describe a computer program that tests whether a pair of cusped hyperbolic manifolds $N_{1}$ and $N_{2}$ has any common Dehn fillings. The program takes $N_{1}$ and $N_{2}$ (or their names) as input, and produces as output a list of pairs of slopes $s_{1}$ on $N_{1}$ and $s_{2}$ on $N_{2}$. For each pair, the manifolds $N_{1}(s_{1})$ and $N_{2}(s_{2})$ are either certified as isometric, or are reported as undistinguished (hence, potentially homeomorphic). As with Section 4.1, our program is not guaranteed to succeed. However, as with that section, all computations apart from length spectra are rigorous. Thus any answer produced by the program is trustworthy provided that systoles are computed correctly to within 1%.
@@
In the second case, SnapPy is unable to rigorously distinguish $V_{1}=\operatorname{vol}(N_{1}(s_{1}))$ from $\operatorname{vol}(N_{2})$. If we had a rigorous proof that $\operatorname{vol}(N_{1}(s_{1}))=\operatorname{vol}(N_{2})$, then as in the first case we would conclude that no fillings of $N_{2}$ can be homeomorphic to $N_{1}(s_{1})$. (Volume coincidences of this form often arise in the setting of arithmetic manifolds. See Proposition 6.2 or a specific instance, and compare [36, Theorem 11.2.3] for a more general statement.) Without a positive lower bound for $\operatorname{vol}(N_{2})-V_{1}$, we cannot use Theorem 3.25 to compute a set of comparison slopes. All we can do in this case is report the approximate equality $\operatorname{vol}(N_{1}(s_{1}))\sim\operatorname{vol}(N_{2})$, and proceed to the next slope.
@@
$\mathcal{V}(V_{1},N_{2})=\left\{s_{2}\text{ a slope on $N_{2}$}\>\bigg{|}\>L(s_{2})<9.97\>\text{ or }\>v\!\left(1-\tfrac{14.77}{L(s_{2})^{2}}\right)>\operatorname{vol}(N_{2})-V_{1}\right\}.$
@@
$\mathcal{T}(s_{1})=\left\{s_{2}\text{ a slope on $N_{2}$}\>\bigg{|}H_{1}(N_{2}(s_{2}))\cong H_{1}(N_{1}(s_{1})),\>\>s_{2}\notin\mathcal{E},\>\>s_{2}\in\mathcal{V}(V_{1},N_{2})\right\}.$
@@
Observe that the manifolds $N_{1}$ and $N_{2}$ play very different roles in Step 5. Thus, to find all common hyperbolic fillings, we need to compare each systole-short slope $s_{2}\in\mathcal{H}(N_{2})$ to the appropriate comparison set of volume-short slopes on $N_{1}$.
@@
We report the sets $\mathcal{U}(N_{i})$ of unidentified slopes on the two manifolds. We report any slopes $s_{i}\in\mathcal{H}(N_{i})$ where the volume of $N_{i}(s_{i})$ could not be distinguished from the volume of the other cusped manifold $N_{j}$. Finally, we report all slope pairs $(s_{1},s_{2})$ where the fillings were confirmed to be isometric and all slope pairs where the fillings could not be distinguished.
@@ 5.2. Results
Among the 12,955 hyperbolic knots with at most 13 crossings, the following 23 knots share a common hyperbolic filling with the figure-8 knot:
@@
Provided that SnapPy’s systole calculations are correct to within 1%, the above is a complete list of knots from the set that are distinct from the figure-8 knot and share a common (nontrivial) Dehn surgery with the figure-8 knot.
@@
[…]
@@
Consider the 1267 hyperbolic knot complements that can be triangulated with fewer than 10 tetrahedra. Assuming that SnapPy computes their systoles to within 1% accuracy, the following hold. There are exactly 61 knots in the set that are distinct from the figure-8 knot and share a common nontrivial Dehn surgery with the figure-8 knot complement. Exactly one of these knots (namely, the $5_{2}$ knot) shares more than one nontrivial Dehn surgery with the figure-8 knot.
@@
Theorems 5.1 and 5.2 are proved by running the program find_common_fillings on the data sets in the two theorems. For the average knot complement $N$ in the two data sets, the search for common fillings between $N$ and the figure–$8$ knot complement takes approximately $5$ seconds.
@@ 6. Volumes of arithmetic manifolds
This section uses several deep but standard results about arithmetic hyperbolic 3-manifolds to establish Proposition 6.2. This proposition is used in the proof of Theorems 5.1 and 5.2. Below, we give a brief summary of the definitions necessary to formulate results about volume. We refer the reader to Maclachlan and Reid [36] for a very thorough treatment of this material. A quick and accessible survey of arithmetic notions appears in Coulson, Goodman, Hodgson, and Neumann [7, Section 4].

### v3

@@ 4.2. Results
The cosmetic surgery conjecture, Conjecture 1.2, holds for the 59,107 one-cusped manifolds in the SnapPy census.
@@
For each manifold $N$ in the census, the procedure runs to completion. Each time, the set of unidentified slopes is $\mathcal{U}(N)=\emptyset$. (This is a consequence of Dunfield’s prior work [13, Theorem 1.2].) The procedure successfully distinguishes all exceptional pairs in $\mathcal{E}\times\mathcal{E}$ and all hyperbolic pairs in $\mathcal{H}^{p}\times\mathcal{T}^{p}$, apart from exactly 7 slope pairs that are described in detail below. In each case, the two fillings of $N$ produce oppositely oriented copies of the same closed 3-manifold $M$, where $M$ is chiral. Thus these fillings do not contradict Conjecture 1.2, which concerns purely cosmetic fillings.
@@
  • $N=\texttt{m172}$, with $s=(0,1)$ and $t=(1,1)$. This is a well-known example, discovered by Bleiler, Hodgson, and Weeks [4]. As described in detail in [4, Section 3], the two fillings $N(s)$ and $N(t)$ are oppositely oriented copies of the lens space $L(49,18)$. This lens space is chiral, because $49$ does not divide $18^{2}+1$.
@@
  • $N=\texttt{m207}$ with $s=(0,1)$ and $t=(1,0)$. Here, $N(s)$ and $N(t)$ are oppositely oriented copies of L(3,1)#L(3,1)L(3,1)\#L(3,1). We check using Regina that each connected summand is $L(3,1)$ rather than the mirror image $\overline{L(3,1)}$. Since $L(3,1)$ is chiral, so is L(3,1)#L(3,1)L(3,1)\#L(3,1).
  • $N=\texttt{t12043}$ with s=(1,1)s=(1,-1) and $t=(1,1)$. Here, $N(s)$ and $N(t)$ are graph manifolds with two Seifert fibered pieces. The first piece, SFS [D: (2,1) (2,1)], is Seifert fibered over a disk with two singular fibers of slope $1/2$. The second piece, SFS [M/n2: (2,1)], is Seifert fibered over a Möbius band with one singular fiber of slope $1/2$. Both of the Seifert pieces here are amphicheiral. However, the gluing matrix $m=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$ does not respect the symmetries of the pieces, so the resulting graph manifold is chiral.
  • $N=\texttt{t12045}$ with s=(1,1)s=(1,-1) and $t=(1,1)$. Here, $N(s)$ and $N(t)$ are graph manifolds with three Seifert fibered pieces: SFS [D: (2,1) (2,1)] and SFS [A: (2,1)] and SFS [D: (2,1) (2,1)]. The middle piece is Seifert fibered over the annulus, with one singular fiber of slope $1/2$. As in the previous bullet, the Seifert pieces are amphicheiral, but the gluing matrices $m=\left[\begin{smallmatrix}1&\text{-}2\\ 0&\text{-}1\end{smallmatrix}\right]$ and $n=\left[\begin{smallmatrix}1&2\\ 1&1\end{smallmatrix}\right]$ do not respect the symmetries of the pieces.
  • $N=\texttt{t12050}$ with s=(1,1)s=(1,-1) and $t=(1,1)$. Here, $N(s)$ and $N(t)$ are graph manifolds with two Seifert fibered pieces: SFS [D: (2,1) (2,1)] and SFS [D: (4,1) (4,1)]. As in the previous bullet, the Seifert pieces are amphicheiral, but the gluing matrix $m=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$ does not respect the symmetries of the pieces.
@@
  • $N=\texttt{m135}$ with $s=(1,3)$ and $t=(3,1)$. Here, SnapPy was unable to compute verified hyperbolic invariants of $N(s)$, because it could not find any positively oriented spun triangulation.
    This verification problem is addressed by taking covers. The double covers of $N(s)$ and $N(t)$ do admit positively oriented triangulations. By a theorem of Hilden, Lozano, and Montesinos [26, Theorem 2.2], the Chern–Simons invariant $\operatorname{CS}(M)$ is multiplicative under covers, modulo $\pi^{2}/2$. The manifold $M=N(s)$ has six distinct double covers up to isomorphism. Each double cover $\widehat{M}$ satisfies $\operatorname{CS}(\widehat{M})=-0.8224...\mod\pi^{2}/2$, which is verifiably distinct from 0. Thus $CS(M)=-0.4112...\mod\pi^{2}/4$, which is again verifiably distinct from 0. Since $CS(N(s))\neq 0$, we know that $N(s)$ is chiral, hence cannot admit an orientation-preserving isometry to $N(t)$.
  • $N=\texttt{t12051}$ with $s=(1,1)$ and t=(2,1)t=(2,-1). Here, SnapPy successfully computed a verified hyperbolic structure on $M=N(s)$ from a positively oriented spun triangulation. The Chern–Simons invariant $CS(M)$ is indistinguishable from 0, and does not prove that $M$ is chiral.
@@
Since each of the above slope pairs produce distinct oriented 3-manifolds, we conclude that Conjecture 1.2 holds for every $N$ in the census. ∎
@@
When ${\operatorname{sys}}(N)$ is very short, the program has to work somewhat harder, for two reasons. First, the systole itself is harder to compute. Our code performs this task by first drilling a very short curve $\gamma$, computing a hyperbolic structure on $N-\gamma$, and then re-filling $\gamma$; the full computation can take around 10 seconds. Second, the set of short slopes $\mathcal{S}(N)$ computed in (4.1) has size approximately $2\pi\cdot{\operatorname{sys}}(N)^{-1}$. After filtering the hyperbolic slopes by homology, as in Steps 4 and 5, the number of slope pairs to check can be on the order of ${\operatorname{sys}}(N)^{-3/2}$. An extreme example is the manifold $N=\texttt{o9\_00637}$, with ${\operatorname{sys}}(N)\approx 0.001643$. For this $N$, the set $\mathcal{S}(N)$ contained 3698 hyperbolic slopes, which were sorted into 124 homology buckets. Altogether, the program had to compare and distinguish 66,842 hyperbolic slope pairs on $N$, which took approximately 65 seconds.
Moving beyond this extreme example, the program encountered 294 one-cusped manifolds whose systole is shorter than $0.004$. Before the innovation of drilling and filling described in the above paragraph, each of those systoles took over an hour to compute. In the latest run of our code, each of these systoles could be computed and rigorously verified in under 10 seconds.
@@
The shortest systole in the one-cusped SnapPy census is achieved by $N=\texttt{o9\_00639}$, with ${\operatorname{sys}}(N)\approx 0.001502$. This manifold has nonzero Casson invariant, so Test 2.5 rules out purely cosmetic surgeries without using hyperbolic methods.
@@
The chirally cosmetic surgery conjecture for knots in $S^{3}$, Conjecture 1.6, holds on the set of all hyperbolic knots up to 15 crossings.
@@
Theorem 4.6 was proved by running the python program check_mfds_chiral on the set of 313,209 hyperbolic knots up to 15 crossings. For a typical knot complement, the program took 10–15 seconds to compute and rigorously verify the systole. (This is longer than the runtimes in Remark 4.4 because even a simplified triangulation of a typical 15–crossing knot complement often has over 25 tetrahedra.) Then, the program took another several seconds to form the set $\mathcal{S}(N)$ and distinguish the corresponding fillings.
However, there is a small list of 14–crossing knots, for which chirally cosmetic surgeries are particularly hard to distinguish:
@@
These 13 knots were previously identified by Stoimenow [45] as being almost amphicheiral: that is, each $K$ in the list is a mutant of its mirror image. Thus, for $N=S^{3}-K$, the Dehn filling $N(p,q)$ is a genus–$2$ mutant of the mirror image of N(p,q)N(p,-q). Since complex hyperbolic volume is preserved under mutation [42, Theorem 1.1 and Corollary 1.2], it cannot distinguish $N(p,q)$ from N(p,q)N(p,-q).
In practice, we distinguished these fillings using two invariants: rigorously certified length spectra up to length $3.5$, and the first homology of covers up to degree $7$ (see Definition 3.1).
Either of these invariants suffices to distinguish $N(p,q)$ from N(p,q)N(p,-q) for $N$ in the above set. Covers up to degree $5$ are very fast to compute, taking a few seconds at most. This suffices for most but not all pairs $s=(p,q)$ and t=(p,q)t=(p,-q). Covers of larger degree often take hours to compute. In an extreme, the computation of enough cover homology groups to distinguish the Dehn fillings of 14a12813 took over 6 days of wall time making full use of 8 cores, for a total of 48 days and 9 hours of CPU time.
By contrast, the rigorous computation of length spectra up to length $3.0$ or $3.5$ typically takes minutes or tens of minutes per pair of slopes. We needed to search up to this fairly large length because mutant manifolds share the same set of short curves up to some cutoff; compare Futer and Millichap [15]. As a result, distinguishing all the Dehn fillings $N(p,q)$ and N(p,q)N(p,-q) for one of the above manifolds often took several hours. For 14a12813, this method took 10 hours.
@@ 5. Common fillings of different parent manifolds
In this section, we describe how the procedure of Section 4 can be adapted to determine common Dehn fillings of two distinct manifolds. The results of this section were motivated by a question from Kalfagianni and Melby. In their joint work [37], they use our results (specifically Theorem 5.2) to find many new qq–hyperbolic knots in $S^{3}$.
@@ 5.1. The procedure
Here, we describe a computer program that tests whether a pair of cusped hyperbolic manifolds $N_{1}$ and $N_{2}$ has any common Dehn fillings. The program takes $N_{1}$ and $N_{2}$ (or their names) as input, and produces as output a list of pairs of slopes $s_{1}$ on $N_{1}$ and $s_{2}$ on $N_{2}$. For each pair, the manifolds $N_{1}(s_{1})$ and $N_{2}(s_{2})$ are either certified as isometric, or are reported as undistinguished (hence, potentially homeomorphic). As with Section 4.1, our program is not guaranteed to succeed. However, as with that section, all computations are rigorous.
@@
In the second case, SnapPy is unable to rigorously distinguish $V_{1}=\operatorname{vol}(N_{1}(s_{1}))$ from $\operatorname{vol}(N_{2})$. If we had a rigorous proof that $\operatorname{vol}(N_{1}(s_{1}))=\operatorname{vol}(N_{2})$, then as in the first case we would conclude that no fillings of $N_{2}$ can be homeomorphic to $N_{1}(s_{1})$. (Volume coincidences of this form often arise in the setting of arithmetic manifolds. See Proposition 6.1 or a specific instance, and compare [41, Theorem 11.2.3] for a more general statement.) Without a positive lower bound for $\operatorname{vol}(N_{2})-V_{1}$, we cannot use Theorem 3.25 to compute a set of comparison slopes. All we can do in this case is report the approximate equality $\operatorname{vol}(N_{1}(s_{1}))\sim\operatorname{vol}(N_{2})$, and proceed to the next slope.
@@
$\mathcal{V}(V_{1},N_{2})=\left\{s_{2}\text{ a slope on $\partial N_{2}$}\>\bigg|\>L(s_{2})<9.97\>\text{ or }\>v\!\left(1-\tfrac{14.77}{L(s_{2})^{2}}\right)>\operatorname{vol}(N_{2})-V_{1}\right\}.$
@@
$\mathcal{T}(s_{1})=\left\{s_{2}\text{ a slope on $\partial N_{2}$}\>\bigg|H_{1}(N_{2}(s_{2}))\cong H_{1}(N_{1}(s_{1})),\>\>s_{2}\notin\mathcal{E},\>\>s_{2}\in\mathcal{V}(V_{1},N_{2})\right\}.$
@@
Observe that the manifolds $N_{1}$ and $N_{2}$ play very different roles in Step 5. Thus, to find all common hyperbolic fillings, we also need to compare each systole-short slope $s_{2}\in\mathcal{H}(N_{2})$ to the appropriate comparison set of volume-short slopes on $\partial N_{1}$.
@@
We report the sets $\mathcal{U}(N_{i})$ of unidentified slopes on the cusps of the manifolds. We report any slopes $s_{i}\in\mathcal{H}(N_{i})$ where the volume of $N_{i}(s_{i})$ could not be distinguished from the volume of the other cusped manifold $N_{j}$. Finally, we report all slope pairs $(s_{1},s_{2})$ where the fillings were confirmed to be isometric and all slope pairs where the fillings could not be distinguished.
@@ 5.2. Results
Among the 12,955 hyperbolic knots with at most 13 crossings, the following 23 knots are a complete list of knots that share a nontrivial Dehn surgery with the figure-8 knot. All of the common fillings are hyperbolic.
@@
We ran the procedure of Section 5.1 on the set of 12,955 hyperbolic knots with at most 13 crossings. For each $K$ in the set, there were no unidentified slopes and no common exceptional fillings between the figure-8 knot and $K$ (apart from the trivial filling that produces $S^{3}$). This leaves hyperbolic fillings. For each hyperbolic filling of each $K$, apart from two examples ($Q=\texttt{5\_2(8,1)}$ and $R=\texttt{12n242(56,3)}$), our program was able to rigorously decide whether or not the resulting manifold also occurred as a filling of the figure–8 knot.
For the remaining two fillings, namely $Q=\texttt{5\_2(8,1)}$ and $R=\texttt{12n242(56,3)}$, SnapPy could not distinguish their volumes from that of the figure-8 knot complement $N=\texttt{4\_1}$. Indeed, the volumes could not be distinguished because they are equal. By Proposition 6.1, we have an equality of volumes
@@
Among the 1267 hyperbolic knot complements that can be triangulated with fewer than 10 tetrahedra, there are exactly 61 knots in the set that are distinct from the figure-8 knot and share a common nontrivial Dehn surgery with the figure-8 knot complement. Exactly one of these knots (namely, the $5_{2}$ knot) shares more than one nontrivial Dehn surgery with the figure-8 knot.
@@
Theorems 5.1 and 5.2 are proved by running the program find_common_fillings on the data sets in the two theorems. For the average knot complement NN in the two data sets, the search for common fillings between NN and the figure–$8$ knot complement takes approximately 55 seconds.
@@ 6. Volumes of arithmetic manifolds
The main goal of this section is to prove the following result, needed in Theorems 5.1 and 5.2.
Proposition 6.1.
Let $Q$ be the closed 3-manifold produced by $(8,1)$ filling on the $5_{2}$ knot. Let $R$ be the closed 3-manifold produced by $(56,3)$ Dehn filling on the (2,3,7)(-2,3,7) pretzel knot. Then both $Q$ and $R$ are hyperbolic and arithmetic. Furthermore,
$\operatorname{vol}(Q)=\operatorname{vol}(R)=2V_{3}.$
The proof of Proposition 6.1 uses several deep but standard results about arithmetic hyperbolic 3-manifolds. In the next subsection, we give a brief summary of the definitions necessary to formulate results about volume; then, we will apply those arithmetic results to prove Proposition 6.1. We refer the reader to Maclachlan and Reid [41] for a very thorough treatment of this material. A quick and accessible survey of arithmetic notions appears in Coulson, Goodman, Hodgson, and Neumann [8, Section 4].
The reader who prefers to avoid technical arithmetic arguments is invited to skip the rest of this section, and consult Remark 6.2 instead.
Remark 6.2.
In personal correspondence, Craig Hodgson and his student James Clift suggested an alternative method for verifying the equality
$\operatorname{vol}(Q)=\operatorname{vol}(R)=2V_{3}$
using geometric triangulations. Each of the two knot complements before the Dehn filling can be triangulated with three ideal tetrahedra. After Dehn filling, $Q=\texttt{5\_2(8,1)}$ has a spun ideal triangulation consisting of the same three tetrahedra, with the tips of the tetrahedra spinning about the core of the surgery solid torus. The same statement is true of $R=\texttt{12n242(56,3)}$.
Clift has written code that can perform a sequence of 2233 moves, 3322 moves, and 220 moves (see [36, Section 2.39] for definitions) to verify that the three-tetrahedron spun triangulation of $Q$ is scissors congruent to the two-tetrahedron triangulation of the figure–$8$ knot complement. His code uses Snap [8] to follow exact solutions to the gluing equations through the retriangulation moves, with shapes given in the number field $k(\Gamma)={\mathbb{Q}}(e^{\pi i/3})$. The same argument works for $R$. Since the manifolds are scissors congruent, they have equal volumes.
The output of Clift’s code is reproduced in the ancillary files [16]. However, we have chosen to preserve the arithmetic proof of Proposition 6.1 because the code is not publicly available.